# Isotope shift analysis with the $4f^{14}6s^{2}~{}^{1}S_{0}-4f^{13}5d6s^{2}(J=2)$ transition in ytterbium

### v1

@@ Abstract
Measurements of isotope shifts have recently been attracting considrable attentions due to their potential in searching for new forces. We report on the isotope shifts of the $4f^{14}6s^{2}~{}^{1}S_{0}-4f^{13}5d6s^{2}(J=2)$ transition at 431 nm in Yb, based on absolute frequency measurements with an accuracy of $\sim 10$ kHz. Utilizing these data, previous reports for other transitions, and theoretical calculation of electronic structure, analyses of isotope shifts on various aspects are performed, including determination of the hyperfine constants for ${}^{173}$b, assessment of nuclear charge radii, analysis on King plots, and a search for new bosons mediating the force between an electron and a neutron. The analysis motivates further precision measurements on isotope shifts of the narrow-linewidth transitions in ytterbium, not only for the bosonic isotopes but also for the fermionic isotopes.
@@ I Introduction
Isotopes, which differs in the number of neutrons $N$ in the nucleus within the same element, exhibit slightly varied physical properties. Such properties include macroscopic quantities, e.g., phase transition temperatures, reaction rates of chemical reactions, and diffusion constants, as well as nuclear properties specific to nuclei. The nuclear properties manifest themselves particularly well in the difference in the resonant frequency $\nu$ for a transition $\alpha$, which is called an isotope shift. Typically ranging from 100 MHz to 1 GHz, the isotope shift for an optical transition of $\sim 100$ THz is easily observed with laser spectroscopies.
@@
Isotope shifts are conventionally attributed primarily to two sources. One is the mass difference between nuclei changing the reduced mass and correlation of electrons, which is called the mass shift. The other is the variation of the nuclear charge radii distorting the Coulomb potential inside the nucleus, referred to as the field shift. Precisely determining the amount of isotope shifts can potentially serve as a good probe of nuclear structure.
@@
In addition to these two main factors, recent development in analysis of isotope shifts incorporates higher order terms as shown in the following equation [1].
$\nu_{\alpha}^{AA^{\prime}}=$ $F_{\alpha}\langle r^{2}\rangle^{AA^{\prime}}+K_{\alpha}\mu^{AA^{\prime}}+G_{\alpha}^{(4)}\langle r^{4}\rangle^{AA^{\prime}}$ (1)
$+G_{\alpha}^{(2)}[(\langle r^{2}\rangle^{A^{\prime\prime}})^{2}]^{AA^{\prime}}+v_{ne}D_{\alpha}N^{AA^{\prime}}$
@@
Here, the superscripted mass number $A$ specifies the isotope, and $X^{AA^{\prime}}\equiv X^{A}-X^{A^{\prime}}$. ${}_{\alpha}$ denotes the transition. Note that in this paper transitions are labeled with their wavelengths (three-digit number in the unit of nm), and all transitions of ytterbium appearing in this paper are summarized in Table 1. Coefficients $F_{\alpha}$ and $K_{\alpha}$ characterize the field shift and mass shift, respectively. $\langle r^{n}\rangle$ is the $n$th moment of the nuclear charge. $\mu^{A}=1/m^{A}$ is the inverse mass. The third and fourth terms are the quadratic field shift, where $[(\langle r^{2}\rangle^{A^{\prime\prime}})^{2}]^{AA^{\prime}}=(\langle r^{2}\rangle^{AA^{\prime\prime}})^{2}-(\langle r^{2}\rangle^{A^{\prime}A^{\prime\prime}})^{2}$, and the fourth-moment shift, respectively. The fifth term shows the effect of a Yukawa potential $V_{ne}=\hbar cv_{ne}\exp(-m_{\phi}cr/\hbar)/r$ on the electrons besides the Coulomb potential. $v_{ne}=(-1)^{s+1}y_{n}y_{e}/(4\pi\hbar c)$ characterizes a hypothetical new force between a neutron and an electron, where $s$ is the spin of the boson mediating this force, $y_{n,e}$ are coupling constants to a neutron and an electron, and $m_{\phi}$ is the mass of the boson.
@@
The field shift term can be eliminated by a set of isotope shifts for another transition $\beta$. The resulting equation is
$\frac{\nu_{\beta}^{AA^{\prime}}}{\mu^{AA^{\prime}}}$ $=$ $\frac{F_{\beta}}{F_{\alpha}}\frac{\nu_{\alpha}^{AA^{\prime}}}{\mu^{AA^{\prime}}}+K_{\beta,\alpha}+\frac{\langle r^{4}\rangle^{AA^{\prime}}}{\mu^{AA^{\prime}}}G_{\beta,\alpha}^{(4)}$ (2)
$+\frac{[(\langle r^{2}\rangle^{A^{\prime\prime}})^{2}]^{AA^{\prime}}}{\mu^{AA^{\prime}}}G_{\beta,\alpha}^{(2)}+\frac{v_{ne}N^{AA^{\prime}}}{\mu^{AA^{\prime}}}D_{\beta,\alpha},$
@@
where $X_{\beta,\alpha}=X_{\beta}-X_{\alpha}F_{\beta}/F_{\alpha}$ for $X\in\{K,D,G^{(4)},G^{(4)}\}$. The plot between $\nu_{\beta}^{AA^{\prime}}/\mu^{AA^{\prime}}$ and $\nu_{\alpha}^{AA^{\prime}}/\mu^{AA^{\prime}}$ is called King plot and approximately linear [2], as the first two terms in the right hand are major. The breakdown of the linearity is proposed to be used for searches for the new bosons [3, 4, 5], and some advanced analysis techniques are also reported to reduce the effects with large uncertainties [6, 7], along with assessments of which higher order effects are significant [8, 9, 10, 11, 12]. For heavy atoms, normalizing both sides with $\nu_{\alpha}^{AA^{\prime}}/\mu^{AA^{\prime}}$ suppresses the influence of the uncertainty in mass [7]:
$\frac{\nu_{\beta}^{AA^{\prime}}}{\nu_{\alpha}^{AA^{\prime}}}$ $=$ $\frac{F_{\beta}}{F_{\alpha}}+K_{\beta,\alpha}\frac{\mu^{AA^{\prime}}}{\nu_{\alpha}^{AA^{\prime}}}+\frac{\langle r^{4}\rangle^{AA^{\prime}}}{\nu_{\alpha}^{AA^{\prime}}}G_{\beta,\alpha}^{(4)}$ (3)
$+\frac{[(\langle r^{2}\rangle^{xA^{\prime\prime}})^{2}]^{AA^{\prime}}}{\nu_{\alpha}^{AA^{\prime}}}G_{\beta,\alpha}^{(2)}+\frac{v_{ne}N^{AA^{\prime}}}{\nu_{\alpha}^{AA^{\prime}}}D_{\beta,\alpha}$
@@
Among various experimental reports on the isotope shift measurements aimed at searching for new bosons [13, 14, 15, 16], ytterbium (Yb) is one of the most extensively studied atomic species for this purpose, as well as nuclear structure, with the nonlinearity of the King plot [7, 1, 17, 18]. The isotope shifts for the 411, 436, and 467 nm transitions in ${}^{+}$ are measured with $\gtrsim 300$ Hz accuracy. For Yb, precise isotope shift measurements for the 361 and 578 nm transitions are reported. However, these analyses so far only cover bosonic isotopes, because the nuclear spins in fermionic isotopes potentially add extra nonlinearity that can become a background in the search for new bosons. In fact, except for the 578 nm transition in Yb, no precise isotope shift measurements are available for fermionic isotopes; only broad-linewidth transitions utilized for laser cooling have the isotope shift data for the fermionic isotopes. Particularly, regarding a newly observed narrow-linewidth transition at 431 nm [19, 20], the absolute frequency is measured for only a single fermionic isotope [19], and isotope shifts are measured for only bosonic isotopes [20].
@@
Table 1: Transitions in Yb and ${}^{+}$. The numbers listed in $\alpha$ are the wavelengths of the transitions in the unit of nm. The lower-energy state for all transitions are the ground state, which is $6s^{2}~{}^{1}S_{0}$ for Yb and $6s~{}^{2}S_{1/2}$ for ${}^{+}$. The $4f$ orbital is fully filled with 14 electrons unless otherwise specified. $\Gamma$ is the linewidth of the excited state.
Yb ${}^{+}$
$\alpha$ excited state $\Gamma$ $\alpha$ excited state $\Gamma$
@@
In this paper, we report absolute frequencies for the 431 nm transition in ${}^{170,172,173,174,176}$b. This report completes the list of isotope shifts of the 431 transition for all stable isotopes at 10 kHz level. Based on these new measurements and some previously reported isotope shifts for other transitions, hyperfine structure, nuclear charge radii, and King plot are analyzed. The king plot analysis also leads to a search for new bosons mediating additional force between a neutron and an electron. The results provide some insights on the nuclear structure for the fermionic isotopes for ytterbium, motivating the precise isotope shift measurements for other transitions.
@@ II Experimental Measurements
The experimental setup and sequences are previously reported elsewhere [19]. To summarize briefly, $\sim 10^{5}$ atoms at $\sim 30$ $\upmu$K are prepared in a magneto-optical trap (MOT) formed with the 556 nm transition. Atoms are then interrogated by 431 nm probe light at 10 mW power generated by second harmonics generation from the 862 nm light emitted by a titanium-sapphire laser. The frequency of the 862 nm light is stabilized to a ultralow-expansion cavity through a frequency comb and a 1064 nm laser, with all relevant radiofrequency signals referenced to the 10 MHz clock signal supplied from a physical realization of Coordinated Universal Time maintained by the National Metrology Institute of Japan. The 431 nm probe light is retroreflected, so that the atoms are immune to the Doppler shift.
Figure 1: Spectra of the depletion of the MOT due to the 431 nm transition: (a) the $F=7/2$ hyperfine state in ${}^{173}$b with 0.0866(13) mT bias magnetic field applied along the 431 nm probe beam. (b) ${}^{174}$b. Black points show the experimental data, and the red curve is the fit of the datapoints. The blue vertical line is the average frequency obtained from the fit. See the main text for the form of the fit function.
The search for the transition is first performed for each isotope by chirping the frequency of the 431 nm light while it is irradiated to the atoms in the MOT. The chirp rate varies from 200 kHz/s to 2 MHz/s depending on the expected strength of the transition. Once the initial signal of the transition is observed, further precise measurements of the resonant frequency is performed with the frequency of the 431 nm light fixed within a cycle. For the $F=7/2,~{}5/2$, and $3/2$ hyperfine states in ${}^{173}$b, the cycle of activating the 431 nm probe light for 3 ms while the MOT is turned off for a short preiod is repeated over 45 times. For other hyperfine states in ${}^{173}$b and bosonic isotopes, the 431 nm light is continuously irradiated onto atoms for 1 s without turning off the MOT. In both cases, the number of atoms after the irradiation of the 431 nm light normalized by the initial atom number is estimated from the amount of fluorescence from the MOT. Every time atoms are removed and then reloaded to the MOT, the frequency of the 431 nm light is shifted to obtain the spectrum of normalized atom number $R$.
Fig. 1 (a) and (b) show sample spectra for the $F=7/2$ state in ${}^{173}$b and ${}^{174}$b, respectively. For the $F=7/2$ state in ${}^{173}$b, distinct separation of Zeeman sublevels is observed due to a bias magnetic field of 0.0866(13) mT applied only when the MOT is deactivated, whereas the Zeeman sublevels in the ${}^{174}$b spectrum overlap because of the inability of applying arbitrary bias magnetic field with the MOT active. The spectra are fitted with the following equation.
$R=R_{0}-\sum_{i=-F}^{F}A_{|i|}\exp\left[-\frac{\left(\nu-\left(\nu_{0}+i\Delta\nu\right)\right)^{2}}{2\sigma^{2}}\right],$ (4)
@@
The absolute frequency of the transition is calculated by compensating systematic shifts from $\nu_{0}$. Major sources are the AC Stark shift due to the 431 nm probe beam and the Doppler shift. The former is estimated for ${}^{173}$b and the bosonic isotopes separately, due to the potential shift in the relative position between the MOT and the probe beam. The frequency shift per 1 mW probe beam is estimated to be 2.40(1.20) kHz and 0.83(1.41) kHz for ${}^{173}$b and ${}^{174}$b as the representative of the bosonic isotopes, respectively. The Doppler shift is assumed to be zero, because the probe beam is retroreflected. However, the 3.8 kHz offset between the average of the recoil-free doublet and the bottom of the Doppler-broadened dip is added in the uncertainty [19]. For the bosonic isotopes, the frequency shift of 101.8(7.8) kHz due to the MOT is also taken into account, as estimated in Ref. [19].
Obtained absolute frequencies are summarized in Table 2. As for ${}^{168}$b, which cannot be trapped with large enough number in the current setup, the absolute frequency is calculated from that for ${}^{174}$b and the isotope shift reported in Ref. [20]. Differences of the absolute frequencies between bosonic isotopes reported in this work agree well within the uncertainty with the isotope shifts reported in Ref. [20].
For the $F=7/2$, $5/2$, and $3/2$ states in ${}^{173}$b, the Landé’s $g$ factor is measured by applying bias magnetic field along the direction of the incident probe beam (See Appendix for details). The bias magnetic field is calibrated using the Zeeman splitting of the 556 nm transition with a relative uncertainty of 1.5%. The obtained $g$ factors are $g_{F}=0.662(46)$, $0.546(24)$, $0.196(24)$ for $F=7/2$, $5/2$, and $3/2$, respectively. $g_{J}$ obtained by the weighted average of these $g_{F}$ is $1.583(56)$, which is consistent with the value in Ref. [19] within an uncertainty, but is off from theoretically expected $g_{J}=1.5$ and the value in Ref. [20].
Table 2: Absolute frequencies for the 431 transition in Yb isotopes: Regarding ${}^{171}$b and ${}^{173}$b, the measured quantities are the transition frequencies for each $F$ state. The number listed in the line without $F$ represents the center of gravity of all $F$ states.
$A$ F Absolute frequency [kHz] Reference
168 695 170 466 295(17) [20]
170 695 172 220 247(19) this work
171 695 172 739 827(24)

### v2

# Isotope-shift analysis with the $4f^{14}6s^{2}~{}^{1}S_{0}-4f^{13}5d6s^{2}(J=2)$ transition in ytterbium

@@ Abstract
Measurements of isotope shifts have recently been attracting considerable attention due to their potentials in searching for new forces. We report on the isotope shifts of the $4f^{14}6s^{2}~{}^{1}S_{0}-4f^{13}5d6s^{2}(J=2)$ transition at 431 nm in Yb, based on absolute frequency measurements with an accuracy of $\sim 10$ kHz. With these data, the hyperfine constants for 173Yb are determined. To analyze these data further, electronic structure of ytterbium is theoretically calculated. The nuclear charge radii are estimated together with some previously reported isotope-shift data for other transitions. An analysis of the King plot for the $4f^{14}6s^{2}~{}^{1}S_{0}-4f^{13}5d6s^{2}(J=2)$ transition shows a good consistency with other transitions, resulting in a constraint on the existence of new bosons mediating the force between an electron and a neutron. The analysis motivates further precision measurements on isotope shifts of the narrow-linewidth transitions in ytterbium, not only for the even-mass isotopes but also for the odd-mass isotopes.
@@ I Introduction
Isotopes, which differ in the number of neutrons $N$ in the nucleus within the same element, exhibit slightly varying physical properties. Such properties include macroscopic quantities, e.g., phase transition temperatures, reaction rates of chemical reactions, and diffusion constants, as well as nuclear properties specific to nuclei. The nuclear properties manifest themselves particularly well in the difference in the resonant frequency $\nu$ for a transition $\alpha$, which is called an isotope shift. Typically ranging from 100 MHz to 1 GHz, the isotope shift for an optical transition of $\sim 100$ THz is easily observed with laser spectroscopy.
@@
Isotope shifts are conventionally attributed primarily to two sources. One is the mass difference between nuclei changing the reduced mass and correlation of electrons, which is called the mass shift. The other is the variation of the nuclear charge radius perturbing the Coulomb potential inside the nucleus, referred to as the field shift. Precise determination of isotope shifts has served as a good probe of nuclear structure [1, 2].
@@
In addition to these two main factors, recent developments in analysis of isotope shifts incorporates higher order terms as shown in the following equation [3]:
$\nu_{\alpha}^{AA^{\prime}}=$ $F_{\alpha}\delta\langle r^{2}\rangle^{AA^{\prime}}+K_{\alpha}\mu^{AA^{\prime}}+G_{\alpha}^{(4)}\delta\langle r^{4}\rangle^{AA^{\prime}}$ (1)
$+G_{\alpha}^{(2)}[(\delta\langle r^{2}\rangle)^{2}]^{AA^{\prime}}+v_{ne}D_{\alpha}N^{AA^{\prime}}$
@@
Here, the superscripted mass number $A$ specifies the isotope, and $X^{AA^{\prime}}\equiv X^{A}-X^{A^{\prime}}$. The subscripted $\alpha$ denotes the transition. Note that transitions are labeled with their wavelengths (three-digit number in the unit of nm) in this paper. All transitions of ytterbium (Yb) appearing in this paper are summarized in Table 1. Coefficients $F_{\alpha}$ and $K_{\alpha}$ characterize the field shift and mass shift, respectively. $\langle r^{n}\rangle$ is the $n$th moment of the nuclear charge, and $\delta\langle r^{n}\rangle^{AA^{\prime}}$ is its difference between two isotopes. $\mu^{A}=1/m^{A}$ is the inverse mass. The third and fourth terms are the fourth-moment shift and the quadratic field shift, respectively, where $[(\delta\langle r^{2}\rangle)^{2}]^{AA^{\prime}}=(\delta\langle r^{2}\rangle^{AA^{\prime\prime}})^{2}-(\delta\langle r^{2}\rangle^{A^{\prime}A^{\prime\prime}})^{2}$ with $A^{\prime\prime}$ being a reference isotope. The fifth term shows the effect of a Yukawa potential $V_{ne}=\hbar cv_{ne}\exp(-m_{\phi}cr/\hbar)/r$ on the electrons besides the Coulomb potential. $v_{ne}=(-1)^{s+1}y_{n}y_{e}/(4\pi\hbar c)$ characterizes the strength of a hypothetical new force between a neutron and an electron, where $s$ is the spin of the boson mediating this force, $y_{n,e}$ are coupling constants to a neutron and an electron, and $m_{\phi}$ is the mass of the boson.
@@
The field-shift term can be eliminated by a set of isotope shifts for another transition $\beta$. The resulting equation is
$\frac{\nu_{\beta}^{AA^{\prime}}}{\mu^{AA^{\prime}}}$ $=$ $\frac{F_{\beta}}{F_{\alpha}}\frac{\nu_{\alpha}^{AA^{\prime}}}{\mu^{AA^{\prime}}}+K_{\beta,\alpha}+\frac{\delta\langle r^{4}\rangle^{AA^{\prime}}}{\mu^{AA^{\prime}}}G_{\beta,\alpha}^{(4)}$ (2)
$+\frac{[(\delta\langle r^{2}\rangle)^{2}]^{AA^{\prime}}}{\mu^{AA^{\prime}}}G_{\beta,\alpha}^{(2)}+\frac{v_{ne}N^{AA^{\prime}}}{\mu^{AA^{\prime}}}D_{\beta,\alpha},$
@@
where $X_{\beta,\alpha}=X_{\beta}-X_{\alpha}F_{\beta}/F_{\alpha}$ for $X\in\{K,D,G^{(4)},G^{(4)}\}$. The plot between $\nu_{\beta}^{AA^{\prime}}/\mu^{AA^{\prime}}$ and $\nu_{\alpha}^{AA^{\prime}}/\mu^{AA^{\prime}}$ is called the King plot and is approximately linear [4], as the first two terms in the right-hand side of Eq. (2) are major. The breakdown of the linearity is proposed to be used for searches for the new bosons [5, 6, 7], and some advanced analysis techniques are also reported to reduce the effects with large uncertainties [8, 9], along with assessments of which higher order effects are significant [10, 11, 12, 13, 14]. For heavy atoms, normalizing both sides with $\nu_{\alpha}^{AA^{\prime}}/\mu^{AA^{\prime}}$ suppresses the influence of the uncertainty in mass [9]:
$\frac{\nu_{\beta}^{AA^{\prime}}}{\nu_{\alpha}^{AA^{\prime}}}$ $=$ $\frac{F_{\beta}}{F_{\alpha}}+K_{\beta,\alpha}\frac{\mu^{AA^{\prime}}}{\nu_{\alpha}^{AA^{\prime}}}+\frac{\delta\langle r^{4}\rangle^{AA^{\prime}}}{\nu_{\alpha}^{AA^{\prime}}}G_{\beta,\alpha}^{(4)}$ (3)
$+\frac{[(\delta\langle r^{2}\rangle)^{2}]^{AA^{\prime}}}{\nu_{\alpha}^{AA^{\prime}}}G_{\beta,\alpha}^{(2)}+\frac{v_{ne}N^{AA^{\prime}}}{\nu_{\alpha}^{AA^{\prime}}}D_{\beta,\alpha}$
@@
Among various atoms[15, 16, 17, 18, 9, 3, 19, 20], Yb is one of the most extensively studied atomic species on the isotope-shift measurements aimed at searching for new bosons, with the nonlinearity of the King plot [9, 3, 19, 20]. The isotope shifts for the 436-nm transitions in Yb+ [9] and the 361-nm transition in Yb [19] are measured with at best 300-Hz accuracy. The 578-nm transition in Yb [20] and the 411- and 467-nm transitions in Yb+ [21] have $\lesssim 10$ Hz accuracy measurements. However, these analyses so far only cover even-mass isotopes, because extra nonlinearity potentially induced by the nuclear spins in odd-mass isotopes can become a background in the search for new bosons. In fact, except for the 578-nm transition in Yb, no precise isotope-shift measurements are available for odd-mass isotopes; only broad-linewidth transitions utilized for laser cooling have the isotope-shift data for the odd-mass isotopes. Particularly, regarding a newly observed narrow-linewidth transition at 431 nm [22, 23], the absolute frequency is measured for only a single odd-mass isotope [22], and isotope shifts are measured for only even-mass isotopes [23].
@@
Table 1: Transitions in Yb and Yb+. The numbers listed in $\lambda$ are the wavelengths of the transitions in the unit of nm. The lower-energy state for all transitions is the ground state, which is $6s^{2}~{}^{1}S_{0}$ for Yb and $6s~{}^{2}S_{1/2}$ for Yb+. The $4f$ orbital is fully filled with 14 electrons unless otherwise specified. $\Gamma$ is the full natural linewidth of the transition. For some transitions $\Gamma$ varies between isotopes.
Yb Yb+
$\lambda$ excited state $\Gamma/2\pi$ $\lambda$ excited state $\Gamma/2\pi$
@@
In this paper, we report absolute frequencies for the 431-nm transition in 170,172,173,174,176Yb. This report completes the list of isotope shifts of the 431-nm transition for all stable isotopes at 10-kHz level. Based on these new measurements and some previously reported isotope shifts for other transitions, hyperfine structure, nuclear charge radii, and the King plot are analyzed. The King plot analysis also leads to a search for new bosons mediating additional force between a neutron and an electron. The results provide some insights on the nuclear structure for the odd-mass isotopes for Yb, motivating precise isotope-shift measurements for other transitions.
@@ II Experimental Measurements
The experimental setup and sequences are previously reported elsewhere [22]. To summarize briefly, $\sim 10^{5}$ atoms at $\sim 30$ $\upmu$K are prepared in a magneto-optical trap (MOT) formed with the 556-nm transition. Atoms are then interrogated by 431-nm probe light at 10-mW power generated by second harmonic generation from the 862-nm light emitted by a titanium-sapphire laser. The frequency of the 862-nm light is stabilized to a ultralow-expansion cavity through a frequency comb and a 1064-nm laser, with all relevant radiofrequency signals referenced to the 10-MHz clock signal supplied from a physical realization of Coordinated Universal Time maintained by the National Metrology Institute of Japan. The 431-nm probe light is retroreflected, so that the atoms are immune to the Doppler shift.
Figure 1: Spectra of the depletion of the MOT due to the 431-nm transition: (a) the $F=7/2$ hyperfine state in 173Yb with 0.0866(13)-mT bias magnetic field applied along the 431-nm probe beam. (b) 174Yb. Black points show the experimental data, and the red curve is the fit of the datapoints. The blue vertical line is the average frequency obtained from the fit. See the main text for the form of the fit function.
The search for the 431-nm transition is first performed for each isotope by chirping the frequency of the 431-nm light while it is irradiated to the atoms in the MOT. The chirp rate varies from 200 kHz/s to 2 MHz/s depending on the expected strength of the transition. Once the initial signal of the transition is observed, further precise measurement of the resonant frequency is performed with the frequency of the 431-nm light fixed within a cycle. For the $F=7/2,~{}5/2$, and $3/2$ hyperfine states in 173Yb, the cycle of activating the 431-nm probe light for 3 ms while the MOT is turned off for a short period is repeated over 45 times. For other hyperfine states in 173Yb and even-mass isotopes, the 431-nm light is continuously irradiated onto atoms for 1 s without turning off the MOT. In both cases, the number of atoms after the irradiation of the 431-nm light normalized by the initial atom number is estimated from the amount of fluorescence from the MOT. Every time atoms are removed and then reloaded to the MOT, the frequency of the 431-nm light is shifted to obtain the spectrum of normalized atom number $R$.
Fig. 1 (b) and (a) show sample spectra for 174Yb and the $F=7/2$ state in 173Yb, respectively. For the $F=7/2$ state in 173Yb, distinct separation of Zeeman sublevels is observed due to a bias magnetic field of 0.0866(13) mT applied only when the MOT is deactivated, whereas the Zeeman sublevels in the 174Yb spectrum overlap because of the inability of applying arbitrary bias magnetic field with the MOT active. The spectra are fitted with the following equation:
$R=R_{0}-\sum_{m=-F}^{F}A_{|m|}\exp\left[-\frac{\left(\nu-\left(\nu_{0}+m\Delta\nu\right)\right)^{2}}{2\sigma^{2}}\right],$ (4)
@@
The absolute frequency of the transition is calculated by compensating systematic shifts from $\nu_{0}$. Major sources are the ac Stark shift due to the 431-nm probe beam and the Doppler shift. The former is estimated for 173Yb and the even-mass isotopes separately, due to the potential shift in the relative position between the MOT and the probe beam. The frequency shift per 1-mW probe beam is estimated to be 2.40(1.20) and 0.83(1.41) kHz for 173Yb and 174Yb as the representative of the even-mass isotopes, respectively.
The first-order Doppler shift is canceled out by the retroreflection of the probe beam. The second-order Doppler shift for 30 $\upmu$K is negligible at the level of 1 kHz uncertainty. In the spectroscopy of a narrow-linewidth transition, a doublet of small recoil-free peaks is observable in the Doppler-broadened dip [24, 25]. In theory, the average frequency of this doublet and the bottom of the dip are at the same frequency, but in practice, these two frequencies do not agree. A 3.8 kHz disagreement is observed in our measurement [22], and this is added to the uncertainty . For the even-mass isotopes, the frequency shift of 101.8(7.8) kHz induced by the ac Stark shift due to the MOT beam is also taken into account, as estimated in Ref. [22]. All other systematic shifts that are carefully measured in the $<1$ Hz level precision spectroscopies, such as second order Zeeman shift [26], Stark shift by ambient fields, collisional shift, blackbody radiation shift, and uncertainty of time base, are well below 1 kHz, and thus their contribution to the uncertainties at this level of accuracy is negligible.
Obtained absolute frequencies are summarized in Table 2. As for 168Yb, which cannot be trapped with large enough number in the current setup, the absolute frequency is calculated from that for 174Yb and the isotope shift reported in Ref. [23]. Differences of the absolute frequencies between even-mass isotopes reported in this paper agree well within the uncertainty with the isotope shifts reported in Ref. [23].
For the $F=7/2$, $5/2$, and $3/2$ states in 173Yb, the Landé $g$ factor is measured by applying bias magnetic field along the direction of the incident probe beam (see Appendix A for details). The bias magnetic field is calibrated using the Zeeman splitting of the 556 nm transition with a relative uncertainty of 1.5%. The obtained $g$ factors are $g_{F}=0.662(46)$, $0.546(24)$, and $0.196(24)$ for $F=7/2$, $5/2$, and $3/2$, respectively. $g_{J}$ obtained by the weighted average of these $g_{F}$ is $1.583(56)$, which is consistent with the value in Ref. [22] within an uncertainty, but is off from theoretically expected $g_{J}=1.5$ and the value in Ref. [23].
Table 2: Absolute frequencies and isotope shift $\nu_{431}^{A176}$ for the 431-nm transition in Yb isotopes: regarding 171Yb and 173Yb, the measured quantities are the transition frequencies for each $F$ state. The number listed in the line without $F$ represents the center of gravity of all $F$ states.
$A$ F Absolute frequency [kHz] $\nu_{431}^{A176}$ [kHz] Reference
168 695 170 466 295(17) -5 680 383(25) [23]
170 695 172 220 247(19) -3 926 431(26) this work
171 695 172 739 889(24) -3 406 789(30)